If Z1,Z2 are part of C such that Im(z1) = sqrt(3), |z2| = 2 and z1*z2 = 4(cos( 5pi/6 ) + sin( 5pi/6)i ). Find the values of Z1 and Z2.

Answer by CPhill(2189)

This problem requires using the properties of complex numbers in both rectangular and polar forms, along with the geometric interpretation of multiplication.
Let $z_1 = x_1 + i y_1$ and $z_2 = x_2 + i y_2$.
### 1. Analyze the Given Information
* **$z_1$ in Rectangular Form:** We are given $\operatorname{Im}(z_1) = \sqrt{3}$.
$$z_1 = x_1 + i\sqrt{3}$$
* **$z_2$ in Polar Form:** We are given $|z_2| = 2$.
Let $z_2 = 2(\cos\theta_2 + i\sin\theta_2)$.
* **Product ($z_p$):** $z_p = z_1 z_2 = 4\left(\cos\left(\frac{5\pi}{6}\right) + i\sin\left(\frac{5\pi}{6}\right)\right)$.
### 2. Convert the Product to Rectangular Form
First, find the exact value of the product $z_p$.
$$\cos\left(\frac{5\pi}{6}\right) = -\cos\left(\frac{\pi}{6}\right) = -\frac{\sqrt{3}}{2}$$
$$\sin\left(\frac{5\pi}{6}\right) = \sin\left(\frac{\pi}{6}\right) = \frac{1}{2}$$
$$z_p = 4\left(-\frac{\sqrt{3}}{2} + i\frac{1}{2}\right)$$
$$z_p = -2\sqrt{3} + 2i$$
### 3. Use the Modulus Property of the Product
The modulus of the product is the product of the moduli: $|z_1 z_2| = |z_1| |z_2|$.
From the product's polar form, $|z_1 z_2| = 4$.
We are given $|z_2| = 2$.
$$4 = |z_1| \cdot 2$$
$$|z_1| = 2$$
### 4. Determine $z_1$
We have $z_1 = x_1 + i\sqrt{3}$ and $|z_1| = 2$.
The modulus is calculated as $|z_1| = \sqrt{x_1^2 + y_1^2}$:
$$2 = \sqrt{x_1^2 + (\sqrt{3})^2}$$
$$4 = x_1^2 + 3$$
$$x_1^2 = 1$$
$$x_1 = \pm 1$$
Therefore, there are two possible solutions for $z_1$:
$$\mathbf{z_{1a} = 1 + i\sqrt{3}} \quad \text{or} \quad \mathbf{z_{1b} = -1 + i\sqrt{3}}$$
### 5. Determine $z_2$ (using $z_2 = \frac{z_p}{z_1}$)
We use the relationship $z_2 = \frac{z_p}{z_1}$, where $z_p = -2\sqrt{3} + 2i$.
#### Case A: Using $z_{1a} = 1 + i\sqrt{3}$
To divide, it's simplest to use the polar form for $z_{1a}$.
$|z_{1a}| = 2$.
$$\arg(z_{1a}) = \arctan\left(\frac{\sqrt{3}}{1}\right) = \frac{\pi}{3}$$
$$z_{1a} = 2\left(\cos\left(\frac{\pi}{3}\right) + i\sin\left(\frac{\pi}{3}\right)\right)$$
The argument of the quotient is the difference of the arguments: $\arg(z_2) = \arg(z_p) - \arg(z_{1a})$.
From $z_p = 4\left(\cos\left(\frac{5\pi}{6}\right) + i\sin\left(\frac{5\pi}{6}\right)\right)$, we have $\arg(z_p) = \frac{5\pi}{6}$.
$$\theta_{2a} = \frac{5\pi}{6} - \frac{\pi}{3} = \frac{5\pi}{6} - \frac{2\pi}{6} = \frac{3\pi}{6} = \frac{\pi}{2}$$
Since $|z_2| = 2$:
$$z_{2a} = 2\left(\cos\left(\frac{\pi}{2}\right) + i\sin\left(\frac{\pi}{2}\right)\right) = 2(0 + i\cdot 1)$$
$$\mathbf{z_{2a} = 2i}$$
#### Case B: Using $z_{1b} = -1 + i\sqrt{3}$
Polar form for $z_{1b}$:
$|z_{1b}| = 2$.
$$\arg(z_{1b}) = \pi - \arctan\left(\frac{\sqrt{3}}{1}\right) = \pi - \frac{\pi}{3} = \frac{2\pi}{3}$$
$$z_{1b} = 2\left(\cos\left(\frac{2\pi}{3}\right) + i\sin\left(\frac{2\pi}{3}\right)\right)$$
$$\theta_{2b} = \arg(z_p) - \arg(z_{1b})$$
$$\theta_{2b} = \frac{5\pi}{6} - \frac{2\pi}{3} = \frac{5\pi}{6} - \frac{4\pi}{6} = \frac{\pi}{6}$$
Since $|z_2| = 2$:
$$z_{2b} = 2\left(\cos\left(\frac{\pi}{6}\right) + i\sin\left(\frac{\pi}{6}\right)\right) = 2\left(\frac{\sqrt{3}}{2} + i\frac{1}{2}\right)$$
$$\mathbf{z_{2b} = \sqrt{3} + i}$$
### Final Solution
There are two pairs of complex numbers $(z_1, z_2)$ that satisfy the given conditions:
$$\text{Pair 1: } \mathbf{z_1 = 1 + i\sqrt{3}} \quad \text{and} \quad \mathbf{z_2 = 2i}$$
$$\text{Pair 2: } \mathbf{z_1 = -1 + i\sqrt{3}} \quad \text{and} \quad \mathbf{z_2 = \sqrt{3} + i}$$